Word problem:
Mrs. Carter owns a shop specializing in tea. She combines 1 lb. Darjeeling with 3 lbs. Earl Grey to get a 4lb mixture of 25% Darjeeling. SHe chooses to change the mixture to 50% Darjeeling. How much Darjeeling should she add to bring the mixture to 50%?
I do not understand how to set this up.

Answer by stanbon(75887)

Mrs. Carter owns a shop specializing in tea.
She combines 1 lb. Darjeeling with 3 lbs. Earl Grey to get a 4lb mixture of 25% Darjeeling.
SHe chooses to change the mixture to 50% Darjeeling. How much Darjeeling should she add to bring the mixture to 50%?
------
Equation:
Dar + Dar = Dar
0.25*4 + x = 0.50(4+x)
Multiply thru by 100 to get:
25*4 + 100x = 50*4 + 50x
50x = 25*4
x = 2 lb (amt. of Darjeeling she must add)
==============================================
Cheers,
Stan H.
================